Answer all these questions. Thank you!

1.
Problem 5.12 (Excel Video) Karen White is saving to buy a house in five years. She plans to put 20 percent down at that time, and she believes that she will need $29,000 for the down payment. If Karen can invest in a fund that pays 9.40 percent annual interest, compounded quarterly, how much will she have to invest today to have enough money for the down payment? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest penny.) Excel Template (Note: This template includes the problem statement as it appears in your textbook. The problem assigned to you here may have different values. When using this template, copy the problem statement from this screen for easy reference to the values youve been given here, and be sure to update any values that may have been pre-entered in the template based on the textbook version of the problem.) Amount to be invested today
2.
Problem 5.16 You invest $190 in a mutual fund today that pays 7.70 percent interest annually. How long will it take to double your money? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to o decimal place, e.g.545) Number of years
3.

You decide to take advantage of the current online dating craze and start your own Web site. You know that you have 550 people who will sign up immediately and, through a careful marketing research and analysis, determine that membership can grow by 29 percent in the first two years, 22 percent in year 3, and 18 percent in Year 4. How many members do you expect to have at the end of four years? (Round intermediate calculations to 6 decimal places, in all cases round your final answer to the nearest whole number.) Excel Template (Note: This template includes the problem statement as it appears in your textbook. The problem assigned to you here may have different values. When using this template, copy the problem statement from this screen for easy reference to the values youve been given here, and be sure to update any values that may have been pre-entered in the template based on the textbook version of the problem.) Members at the end of four years
4.
Debra plans to invest some money so that she has $4,900 at the end of three years. Determine how much should she invest today given the following choices: (Do not round intermediate calculations and round your final answer to the nearest penny.) a. 4.2 percent compounded daily. Amount required to be invested b. 4.9 percent compounded monthly. Amount required to be invested c. 5.2 percent compounded quarterly. Amount required to be invested d. 5.4 percent compounded annually. Amount required to be invested
5.
Blossom Corp had sales of $382,000 in 2017. If rate at which the co intermediate calculations to 4 decimal places, in all cases round 8.72%.) management expects its sales to be $476,450 in 4 years, what is the mpanys sales are expe cted to grow? (If you solve this problem with algebra round your final answer to 2 decimal places, e.g. Growth rate
Answer #1

1)Here FV = 29000$
R = rate of interest = 9.4%/4 = 2.35%
N = 5*4 =20
PV = FV/(1+r)^n
PV = 29000/(1+2.35%)^20
PV = 29000/(1.0235)^20
=29000/1.591318
= 18223.89$
Thus she need to invest $ 18223.89 today

2) Here PV = 190$
R = 7.7%
FV = 190*2 =380$
n = ?
FV =PV(1+r)^n
380 =190(1.077)^n
(1.077)^n =2
assuming n = 9
(1.077)^9 =1.9496
assuming n = 10
(1.077)^10 = 2.0996
using intrapolation we get
n = 9.336 years

N Figure
9 1.9496
10 2.0996
1 0.15
0.0504

=0.0504/0.15 = 0.336

3) table showing number of member at year 4

Year Growth(in%) Number of member
0 550
1 29% 710
2 29% 915
3 22% 1117
4 18% 1318
Thus number of member at end of 4 year is 1318

4)

Here FV = 4900$
A) R = 4.2%/365 = 0.0115%
n=3*365 =1095
PV = FV/(1+r)^n
=4900/(1.000115)^1095
=4900/1.134189
=4320.268$
B) R = 4.9%/12 = 0.4083%
N= 3*12 =36
PV = FV/(1+r)^n
=4900/(1.004083)^36
=4900/1.157993
=4231.457$
C) R = 5.2%/4 = 1.3%
n= 3*4 =12
PV = FV/(1+r)^n
=4900/(1.013)^12
=4900/1.167652
=4196.457$
D)R = 5.4%
n= 3
PV = FV/(1+r)^n
=4900/(1.054)^3
=4900/1.170905
=4184.796$

5) Here PV = 382000
FV = 476450
N = 4 years
FV = PV(1+r)^n
=476450=382000(1+r)^4
(1+r)^4 =1.24725
(1+r)^4*1/4 = 1.12475^1/4
(1+r) = 1.029826
r=0.029826
ie 2.9826%